EXPLORING SHOCK CHEMISTRY IN ORION-KL WITH MID-J MOLECULAR TRANSITIONS

\underline{JULIE K. ANDERSON} and LUCY M. ZIURYS, Department of Chemistry and Biochemistry, The University of Arizona, Tucson, AZ 85719.

New receiver technology can allow us to probe high energy shock chemistry using ground-based telescopes. We have mapped HCN (J = 8 → 7) and HCO^{+} (J = 8 → 7) emission across the 1′ x 1′ Orion-KL region using the ALMA Band 9 receiver at the Submillimeter Telescope of the Arizona Radio Observatory. A map of the SiO (J = 16 → 15) emission is also currently underway. The mid-J HCN and HCO^{+} emission have very broad line widths that vary with position, which suggest an association with the high velocity globules produced by the explosive outflow in this region. The emission points to densities greater than 4x10^{8} cm^{−3} in the shock material. Additionally, the central velocity of the molecular lines changes across the nebula, eluding to interesting globular structures containing the molecular material.